Resolved

A resolution for an open issue

Unscheduled downtime - power issue at primary data centre

Date: 
Thu, 12/01/2017 - 10:40

We are currently experiencing an issue with connectivity at our primary data centre resulting in the unavailability of customer servers.

UPDATE 10:52 - We have been informed that there is a power issue at our primary data centre which explains the unscheduled downtime. Engineers at the data centre are looking into this as a matter of urgency and we'll update this page as soon as we have further information.

Storage issue affecting some virtual machines

Date: 
Sat, 28/11/2015 - 20:05 - 21:30

We are currently experiencing a storage issue that has caused a small number of customer virtual machines to become unavailable. One of the disk groups on our storage network has become full unexpectedly this evening, causing the storage to go offline for some customers.

We are currently working to resolve the issue by bringing the affected disks back online and restarting the affected virtual machines.

We will also be investigating the underlying issue and improving our monitoring in due course to ensure that this situation will not reoccur.

Network Downtime

Date: 
Tue, 18/08/2015 - 10:55 - 11:15

We experienced an Internet connectivity issue at our primary data centre, which affected most customer services for 20 minutes. We have confirmed this was caused by an issue with our upstream provider's network and are awaiting further information.

The network has now been stable for over 30 minutes and we are not expecting any further disruption. We apologise for any inconvenience caused.

Problems with connectivity to main datacentre

Date: 
Mon, 28/10/2013 - 13:20 - 13:30

We experienced a Network outage at our main Data centre this afternoon around 13:20pm (GMT). The issue occurred on the backbone network of the Data centre and was fixed within 10 minutes of going down. This has affected numerous other customers using this data centre and is not an exclusive Network EQ issue. We shall be receiving an RCA (Root Cause Analysis) shortly from the data centre provider and will be reflecting the RCA here as soon as we receive it.

If you are still experiencing connection issues, please let us know asap.

Apologies for any inconvenience caused by this outage.

Problem with several servers offline

Date: 
Fri, 08/03/2013 - 08:52

We are currently experiencing issues with one of our Xen hosts which has caused some servers to become unavailable.

We are working on getting the servers back online as soon as possible.

We apologise for any inconvenience caused.

Update 08/03/13 9:15AM: All servers should now be back online and functioning correctly, we are still investigating the issue and will keep this site up to date with the latest information we have.

Issue with Sending and Receiving email on Cpanel01 (Shared Hosting)

Date: 
Wed, 06/03/2013 - 13:20 - 14:09

We currently have an issue with sending and receiving email on one of our shared hosting servers. We are looking into the issue now and will provide updates here as soon as we can.

We apologise for any inconvenience caused.

Update 14:09: We believe we have now fixed the issue which was to do with sender verification not working correctly. You email should start to be sent and received again in the next five minutes or so. If you do have any issues after this time, please phone our support team on 03333441270

Virtual Server Disk Issue

Date: 
Wed, 06/03/2013 - 02:15 - 08:20

A storage issue is continuing to affect some virtual servers and services at our primary site. Our engineers are working on the issue.

We apologise for any inconvenience caused.

Update 06/03/2013 06:05: The issue has been rectified and our engineers are bringing the affected virtual machines back online as fast as possible, but the association between disks and virtual machines is having to be re-created manually so this will take a little time.

Migration From Leading Edge Hosting to Network EQ Infrastructure

Date: 
Wed, 26/12/2012 - 19:24

Any Domains, Websites and Email that has been stated to be moved between the 20th of December and the 26th December has been postponed to the 27th and 28th Dec due to a backend issue with MySQL. We took the decision to not move any more websites until this was resolved and whilst the MySQL issue has been resolved, we will be taking precautionary measures in the moves of the accounts in the next few days.

If you do experience any issues with the account migrations, please contact us immediately on 03333441270.

We apologise for any inconvenience caused.

Planned Maintenance of cPanel Servers

Date: 
Tue, 20/11/2012 - 11:00 - 11:15

We will be restarting the cPanel servers tonight around 11pm to perform several maintenance tasks of the servers.
The outage will last a few minutes but problems connecting to websites may be experienced between 11pm and 11:15pm.

We apologise for any inconvenience this may cause.

Update 23:15pm: The planned maintenance is now complete and all services are running as normal.
We apologise for any inconvenience caused.

Connectivity Loss

Date: 
Fri, 02/11/2012 - 14:54

We are currently experiencing a connectivity issue which is affecting all Network EQ hosting Customers, those who have Leading Edge services and Hosted Exchange services remain unaffected.

Please bear with us whilst we try to resolve the issue.

We apologise for any inconvenience caused.

Update 15:49pm: All connectivity has been back up for about 50 minutes with no further loss of connection. We are looking into the cause of the issue and will be taking preventative steps in the planned maintenance tomorrow night.

Status Update Feed